# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 5 The vector FL equation
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue. Let $u$ be a solution of the vector FL equation (5.1) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$, where $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$. Furthermore, let $\tilde{\eta}$ be an $n\times m_{2}$ matrix solution of the linear system
@@
The following is a particular case of the “superposition property” formulated in Remark 4.4. Let $(\Gamma^{(i)},\tilde{\eta}^{(i)},\Omega^{(i)})$, $i=1,2$, be data determining solutions ${u^{\prime}}^{(i)}$ of the vector FL equation via the vectorial Darboux transformation, using the same seed solution $u$. Let
@@
Inserting the plane wave ansatz
@@
which requires that the scalars $u^{\dagger}\boldsymbol{\sigma}u$ and $u^{\dagger}\boldsymbol{\sigma}\boldsymbol{\alpha}\,u$ are constant. This is guaranteed is $\boldsymbol{\alpha}$ commutes with $\boldsymbol{\sigma}$.
With certain restrictions, in Sections 5.1 and 6, this plane wave solution will be taken as a seed solution $u$ in the above Darboux transformation for the vector FL equation. The assumptions $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$ in Theorem 5.1 now read
@@ 6 Solutions of a two-component vector FL equation
In this section, we elaborate in detail the set of solutions of the two-component vector FL equation, i.e, (5.1) with $m_{2}=2$, generated via the Darboux transformation obtained in the preceding section. In order to somewhat reduce the complexity, we restrict our considerations to the case where $\boldsymbol{\sigma}=I_{2}$ and $\boldsymbol{\alpha}=\mathrm{diag}(\alpha_{1},\alpha_{2})$, with $\alpha_{\mu}\in\mathbb{R}$, $\mu=1,2$. Then we can choose
@@
which satisfies $(A_{1}e^{\mathrm{i}\,\varphi_{1}},A_{2}e^{\mathrm{i}\,\varphi_{2}})^{T}=\boldsymbol{U}\,(\sqrt{A^{\dagger}\boldsymbol{\sigma}A},0)^{T}$ and the $\boldsymbol{\sigma}$-unitarity condition $\boldsymbol{U}^{\dagger}\boldsymbol{\sigma}\boldsymbol{U}=\boldsymbol{\sigma}$. If $\sigma=-1$, so that $\boldsymbol{\sigma}$ is indefinite, we have to restrict $A$ such that $|A_{2}|<|A_{1}|$. $\Box$
6.1 The special case $\alpha_{1}=\alpha_{2}$
@@
The linear system then reads
$\chi_{1xx}-\mathrm{i}\,\alpha\,\chi_{1x}-\frac{1}{4}\big{(}\Gamma^{-1}-2\mathrm{i}\,\alpha^{2}\beta\,I_{n}\big{)}\,\Gamma^{-1}\,\chi_{1}=0\,,\qquad\chi_{2x}-\Big{(}\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\alpha\,I_{n}\Big{)}\,\chi_{2}=0\,$,
$\chi_{1t}+\frac{\mathrm{i}\,}{\alpha}\Gamma\,\chi_{1x}+\frac{1}{2}\big{(}\Gamma-\mathrm{i}\,\beta\,I_{n}\big{)}\,\chi_{1}=0\,,\qquad\chi_{2t}-\Big{(}\frac{1}{2}\Gamma+\mathrm{i}\,\beta\,I_{n}\Big{)}\,\chi_{2}=0\,.$
It is (in special cases not necessarily completely) solved by
$\chi_{1}=e^{\frac{\mathrm{i}\,}{2}\varphi}\,\sum_{\kappa}e^{R_{\kappa}(I_{n}x-\frac{\mathrm{i}\,}{\alpha}\Gamma\,t)}\,c_{\kappa}\,,\qquad\chi_{2}=e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,c_{0}\,$, (6.12)
where $c_{0}$ and $c_{\kappa}$ are constant $n$-component column vectors, $\{R_{\kappa}\}$ is a set of different roots of
@@
$\varphi=\alpha\,x+\beta\,t=\alpha\,x+(2\|A\|^{2}-1/\alpha)\,t\,.$
After determining $\Omega$, finally,
@@
In this case, $R_{\kappa}$ is diagonal, $R_{\kappa}=\mathrm{diag}(r_{1\kappa},\ldots,r_{n\kappa})$.
@@
$\Omega_{t}$ $=$ $-\frac{\mathrm{i}\,}{2\alpha^{2}\|A\|^{4}}\,\Xi\,\Xi^{\dagger}\,\Gamma^{\dagger}-\frac{\mathrm{i}\,}{\alpha\,\|A\|}\Xi\,\chi_{1}^{\dagger}+\frac{1}{2}(\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger})\,.$
6.1.1 The case $n=1$
For $n=1$, the roots of the quadratic equation (6.13) are $r_{\pm}=\pm w/(2\gamma)$ with
@@ 6.1.1 The case $n=1$
Let $\mathrm{Re}(\gamma)\neq 0$ and also $w\neq 0$. (6.12) yields the general solution of the linear system,
$\chi_{1}=e^{\frac{1}{2}\mathrm{i}\,\varphi}\,\big{(}c_{1}\,e^{\rho}+c_{2}\,e^{-\rho}\big{)}\,,\qquad\chi_{2}=c_{0}\,e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma t)}\,$,
@@
The above expressions also cover the alternative case $c_{1}=0$ (and $c_{2}\neq 0$) if we allow for $w$ also the negative of the square root in (6.23). If $\mathrm{Re}\big{(}\gamma\,(w^{\ast}-1)\big{)}=0$, then both components of $u^{\prime}$ are functions of $t$ only, multiplied by a phase factor with an exponent linear in $x$. See Appendices Appendix A: Special examples from the class of solutions in Example 6.2 and Appendix B: An almost trivial class of solutions of the two-component vector FL equation for this case. If $\mathrm{Re}\big{(}\gamma\,(w^{\ast}-1)\big{)}\neq 0$, then the wave, described by $u^{\prime}$, travels with constant velocity
@@
As a consequence, $u_{1}^{\prime}$ involves a dark soliton if $\alpha\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\gamma\,(w^{\ast}-1))<0$. If the maximum of the absolute value in the expression for $|u_{1}^{\prime}|^{2}$ is smaller than 1, we have a single dark soliton. An example is shown in Fig. 3.
Figure 3: Plots of the absolute value of the first, and the real part (the imaginary part is similar) of the second component for an $n=1$ solution from the class in Example 6.2. The parameters are $\alpha=\|A\|=2$, $\gamma=1+\mathrm{i}\,$, $c_{0}=c_{1}=1$, $c_{2}=0$, but we replaced $w$ by w-w. The second plot shows the “beating”. The velocity is $v\approx 0.3$, the beating period $T^{\prime}\approx 1$.
@@
In each case, $u_{2}^{\prime}$ describes a bright soliton (on vanishing background). $\Box$
@@
Let $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$. Then the two roots of the quadratic equation are
@@
$\chi_{1}=e^{\frac{\mathrm{i}\,}{2}\varphi}\,e^{r_{\pm}(x-\frac{k}{\alpha}\,t)}\,c_{1}\,,\qquad\chi_{2}=0\,$,
where $c_{1}\neq 0$. The Lyapunov equation then reduces to the constraint
$k\,\big{|}r_{\pm}-\frac{\mathrm{i}\,}{2}(\alpha+\frac{1}{k})\big{|}^{2}=\alpha^{2}\|A\|^{2}\,$,
which requires $k>0$. The differential equations for $\Omega$ read
$\Omega_{x}=\frac{|c_{1}|^{2}}{2\,k^{2}}\big{(}1+2\,\mathrm{i}\,\,k\,(r_{\pm}+\mathrm{i}\,\frac{\alpha}{2})\big{)}\,e^{2\theta_{\pm}}\,,\qquad\Omega_{t}=-\frac{k}{\alpha}\,\Omega_{x}\,$,
@@
$\theta_{\pm}=r_{\pm}\,\big{(}x-\frac{k}{\alpha}\,t\big{)}\,.$
(i) $r_{\pm}\neq 0$ real. The constraint is then automatically satisfied and we obtain
$\Omega=\frac{\mathrm{i}\,\,|c_{1}|^{2}\,(1+\mathrm{i}\,k\,(2r_{\pm}+\mathrm{i}\,\alpha))}{2\,k^{2}\,(\alpha+\mathrm{i}\,\,(2r_{\pm}+\mathrm{i}\,\alpha))}\,e^{2\theta_{\pm}}+c\,$,
@@
$u_{\mu}^{\prime}$ $=$ $A_{\mu}\,e^{\mathrm{i}\,\big{(}\alpha x+(2\|A\|^{2}-\alpha^{-1})t\big{)}}\,\Big{(}1$
$+\frac{k\,|c_{1}|^{2}\big{(}k\,(2r_{\pm}+\mathrm{i}\,\alpha)-\mathrm{i}\,(1+2\,\alpha\,k)\big{)}\,(\alpha+\mathrm{i}\,(2r_{\pm}+\mathrm{i}\,\alpha))\,e^{2\theta_{\pm}}}{\alpha\|A\|^{2}\big{(}\mathrm{i}\,|c_{1}|^{2}(1+\mathrm{i}\,k\,(2r_{\pm}+\mathrm{i}\,\alpha))\,e^{2\theta_{\pm}}+2\,k^{2}\,c\,(\alpha+\mathrm{i}\,(2r_{\pm}+\mathrm{i}\,\alpha))\big{)}}\Big{)}\,,\quad\mu=1,2\,.$
The components correspond to a solution of the scalar FL equation.
(ii) $r_{\pm}=0$. Then we have
$k=2\,\|A\|^{2}-\alpha^{-1}\pm 2\sqrt{\|A\|^{2}(\|A\|^{2}-\alpha^{-1})}\,.$
This is real if and only if $\|A\|^{2}\geq\alpha^{-1}$. In case of the plus sign, $k$ is then positive. In case of the minus sign, $k$ is only positive if $\alpha<0$. Furthermore, we have $\theta_{\pm}=0$, and thus
@@
$u^{\prime}$ $=$ $e^{\mathrm{i}\,\big{(}\alpha x+(2\,\|A\|^{2}-\alpha^{-1})t\big{)}}\,\Big{(}1-\frac{\mathrm{i}\,|c_{1}|^{2}}{2\alpha\,\|A\|^{2}\,\Omega}\big{(}\alpha+\frac{1}{k}\big{)}\Big{)}\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)$ (6.30)
$=$ $e^{\mathrm{i}\,\big{(}\alpha x+(2\,\|A\|^{2}-\alpha^{-1})t\big{)}}\,\Big{(}1-\frac{\mathrm{i}\,k\,(1+k\alpha)}{\|A\|^{2}\,\big{(}(1-k\alpha)\,(\alpha\,x-k\,t)+\mathrm{i}\,k\alpha\big{)}}\Big{)}\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)\,$, (6.33)
which represents a soliton rationally concentrated along the line $x=(k/\alpha)\,t$, where $|u^{\prime}_{\mu}|=|A_{\mu}|\,|1-(\alpha^{-1}+k)/\|A\|^{2}|$. This shows that we can also get dark solitons in this way. An example is displayed in Fig. 5. Also in this case, the components only differ in their amplitudes and thus correspond to a solution of the scalar FL equation.
@@
(iii) $r_{\pm}\neq 0$ imaginary. Then the above constraint from the Lyapunov equation is not compatible with the expressions for $r_{\pm}$. $\Box$
In case of a double root, i.e., $r_{+}=r_{-}$, (6.12) does not yield the complete solution of the linear system. In this case, we have to find the missing solution. This is done in the next example.
@@
We have a double root $r=0$ if

### v2

@@ 5 The vector FL equation
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue (spectrum condition). Let $u$ be a solution of the vector FL equation (5.1) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$, where $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$. Furthermore, let $\tilde{\eta}$ be an $n\times m_{2}$ matrix solution of the linear system
@@
Remark 5.3.
Using a familiar formula for the determinant of a $2\times 2$ block matrix, the new solution (5.6) can be expressed as follows in terms of a quotient of determinants,
$u^{\prime}_{\mu}$ $=$ $u_{\mu}-\frac{\mathrm{i}}{(u_{x}^{\dagger}\boldsymbol{\sigma}u)\,\det(\Omega)}\,\det\left(\begin{array}[]{cc}\Omega&\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u\\ \boldsymbol{\sigma}\,\tilde{\eta}_{\mu}^{\dagger}&0\end{array}\right)$
$=$ $\frac{1}{\det(\Omega)}\,\det\left(\begin{array}[]{cc}\Omega&\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u\\ \mathrm{i}\,(u_{x}^{\dagger}\boldsymbol{\sigma}u)^{-1}\boldsymbol{\sigma}\,\tilde{\eta}_{\mu}^{\dagger}&u_{\mu}\end{array}\right)\qquad\quad\mu=1,\ldots,m_{2}\,.$
In the cited references, where solutions have been generated via the classical (iterative) Darboux transformation method, they are mostly presented in such a form, with a specialized matrix $\Omega$. $\Box$
Remark 5.4.
@@
We specialize the important “superposition property”, formulated in Remark 4.4, to the case of the vector FL equation. Let $(\Gamma^{(i)},\tilde{\eta}^{(i)},\Omega^{(i)})$, $i=1,2$, be data determining solutions ${u^{\prime}}^{(i)}$ of the vector FL equation via the vectorial Darboux transformation, using the same seed solution $u$. Let
@@
In the following subsection and the next section we concentrate on a plane wave seed solution in the above Darboux transformation for the vector FL equation. Inserting the plane wave ansatz
@@
which requires that the scalars $u^{\dagger}\boldsymbol{\sigma}u$ and $u^{\dagger}\boldsymbol{\sigma}\boldsymbol{\alpha}\,u$ are constant. This is guaranteed is $\boldsymbol{\alpha}$ commutes with $\boldsymbol{\sigma}$. The assumptions $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$ in Theorem 5.1 now read
@@ 6 Solutions of a two-component vector FL equation
In this section, we elaborate the set of solutions of the two-component vector FL equation, i.e, (5.1) with $m_{2}=2$, generated via the Darboux transformation obtained in the preceding section. In order to somewhat reduce the complexity, we restrict our considerations to the case where $\boldsymbol{\sigma}=I_{2}$ and $\boldsymbol{\alpha}=\mathrm{diag}(\alpha_{1},\alpha_{2})$, with $\alpha_{\mu}\in\mathbb{R}$, $\mu=1,2$. Then we can choose
@@
The corresponding analysis is different, depending on whether the spectrum condition for the matrix $\Gamma$ in the Lyapunov equation holds or not. It is necessary to distinguish several cases. Moreover, for special choices of $\Gamma$, the linear system, which here is equivalent to a constant coefficient linear system, “degenerates” and admits quasi-rational solutions. In Sections 6.1 and 6.2, we will organize the various possibilities into a collection of examples.
@@
which satisfies $(A_{1}e^{\mathrm{i}\,\varphi_{1}},A_{2}e^{\mathrm{i}\,\varphi_{2}})^{T}=\boldsymbol{U}\,(\sqrt{A^{\dagger}\boldsymbol{\sigma}A},0)^{T}$ and the $\boldsymbol{\sigma}$-unitarity condition $\boldsymbol{U}^{\dagger}\boldsymbol{\sigma}\boldsymbol{U}=\boldsymbol{\sigma}$. If $\sigma=-1$, so that $\boldsymbol{\sigma}$ is indefinite, we have to restrict $A$ such that $|A_{1}|\neq|A_{2}|$. $\Box$
Remark 6.2.
The form of the two-component vector FL equation in [7, 16] and corresponding solutions are obtained from ours via $x\mapsto 2x$ and $t\mapsto-t/2$. In [10, 11] the two-component vector FL equation is taken in the form
$\mathrm{i}\,D_{\tau}\frac{\partial v_{1}}{\partial\xi}+\frac{\sigma}{2}\frac{\partial^{2}v_{1}}{\partial\tau^{2}}+(2|v_{1}|^{2}+|v_{2}|^{2})\,D_{\tau}v_{1}+v_{1}v_{2}^{\ast}D_{\tau}v_{2}=0\,$,
$\mathrm{i}\,D_{\tau}\frac{\partial v_{2}}{\partial\xi}+\frac{\sigma}{2}\frac{\partial^{2}v_{2}}{\partial\tau^{2}}+(2|v_{2}|^{2}+|v_{1}|^{2})\,D_{\tau}v_{2}+v_{2}v_{1}^{\ast}D_{\tau}v_{1}=0\,$,
where $D_{\tau}=1+\mathrm{i}\,\epsilon\,\partial/\partial\tau$, $\xi$ and $\tau$ are independent variables, and $\epsilon\neq 0$ and $\sigma$ are real constants. Then
$v_{j}(\tau,\xi)=e^{\mathrm{i}\,(\epsilon^{-1}\tau+\sigma\epsilon^{-2}\xi)}\,u_{j}(x,t)\qquad j=1,2\,$,
with $x=-\frac{1}{4}\sigma\epsilon^{-4}\,(\sigma\,\xi+\epsilon\,\tau)$ and $t=\xi$, maps (solutions of) the latter system to (solutions of) our form of the two-component vector FL equation, and vice versa (also see [16]). $\Box$
6.1 The special plane wave case $\alpha_{1}=\alpha_{2}$
@@
Writing
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,\tilde{\chi}_{1}\,,\qquad\chi_{2}=e^{\mathrm{i}\,\varphi}\,\tilde{\chi}_{2}\,$,
with
$\varphi=\alpha\,x+\beta\,t=\alpha\,x+(2\|A\|^{2}-1/\alpha)\,t\,$,
the linear system becomes
$\tilde{\chi}_{1xx}-R^{2}\,\tilde{\chi}_{1}=0\,,\quad\tilde{\chi}_{1t}+\frac{\mathrm{i}}{\alpha}\Gamma\,\tilde{\chi}_{1x}=0\,$, (6.12)
where [^6] If the right hand side does not admit a matrix root, we shall simply regard $R^{2}$ as an abbreviation for the right hand side of (6.13).
@@
$\tilde{\chi}_{2x}-\frac{1}{2}\Gamma^{-1}\tilde{\chi}_{2}=0\,,\quad\tilde{\chi}_{2t}-\frac{1}{2}\Gamma\,\tilde{\chi}_{2}=0\,$,
which is solved by
$\tilde{\chi}_{2}=e^{\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,c_{0}\,$,
where $c_{0}$ is a constant $n$-component column vector.
After determining the solution of the linear system and then $\Omega$, finally,
@@
Given an $n=1$ soliton solution, generated with some value of $\gamma$, an “$n$-th order” extension of it is determined by an $n\times n$ (lower triangular) Jordan block with eigenvalue $\gamma$.
@@
$\Omega_{t}$ $=$ $-\frac{\mathrm{i}}{2\alpha^{2}\|A\|^{4}}\,\Xi\,\Xi^{\dagger}\,\Gamma^{\dagger}-\frac{\mathrm{i}}{\alpha\,\|A\|}\Xi\,\chi_{1}^{\dagger}+\frac{1}{2}(\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger})\,.$ (6.23)
If the expression on the right hand side of (6.13) is invertible, then it indeed possesses a matrix root $R$. (6.12) is then solved by
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,\sum_{\kappa}e^{R_{\kappa}(I_{n}x-\frac{\mathrm{i}}{\alpha}\Gamma\,t)}\,c_{\kappa}\,$, (6.24)
where $c_{\kappa}$ are constant $n$-component column vectors and $\{R_{\kappa}\}$ is a set of different roots of (6.13).
If the right hand side of (6.13) is not invertible, so that typically no square root exists (and $R^{2}$ in (6.12) should then be replaced by the right hand side of (6.13)), we have to step back to (6.12) to determine its solution. This is actually an important case, since it leads in particular to rogue wave solutions.
Let us recall that a solution $u^{\prime}$ is called quasi-rational if it has the form $u^{\prime}=e^{\mathrm{i}\,\varphi}\,(v_{1},v_{2})^{T}$ with (in $x$ and $t$) rational expressions $v_{1},v_{2}$. If $u_{1}$ is a solution of the scalar FL equation, then $u=U_{0}\,(u_{1},0)^{T}$, with a constant unitary matrix $U_{0}$ solves the two-component vector FL equation. In this case we say that $u$ is equivalent to the solution $u_{1}$ of the scalar FL equation.
Proposition 6.3.
Any quasi-rational solution of the two-component vector FL equation, obtained by the vectorial Darboux transformation from the special plane wave seed with $\alpha_{1}=\alpha_{2}$, is equivalent to a quasi-rational solution of the scalar FL equation.
Proof.
In order for (6.21) to be quasi-rational, it is necessary that $\chi_{2}=0$. But then the second component of $u^{\prime}$ in the last expression in (6.21) vanishes. ∎
Remark 6.4.
In [8, 6, 12, 17, 13] only the special ($\alpha_{1}=\alpha_{2}$) plane wave background case has been used as a seed solution, applying a generalized Darboux transformation, with a limit procedure that takes different spectral parameters to the same value, in order to obtain higher order solutions, in particular higher order rogue waves. In our framework, the corresponding $n$-th order solution is obtained by choosing $\Gamma$ to be an $n\times n$ Jordan block. Proposition 6.3 indicates that all quasi-rational solutions presented in the mentioned publications are equivalent to quasi-rational solutions of the scalar FL equation, which is indeed the case. This is a useful insight within the vectorial Darboux transformation framework. The publication [7] also contains solutions with the special plane wave background, but steps into the richer case of different wave numbers ($\alpha_{1}\neq\alpha_{2}$, see Section 6.2), though without treating rogue waves. $\Box$
6.1.1 $n=1$
For $n=1$, the roots of the quadratic characteristic equation (6.13) are $r_{\pm}=\pm w/(2\gamma)$ with
@@ 6.1.1 $n=1$
Throughout this subsection, $\chi_{2}$ is always given by
$\chi_{2}=c_{0}\,e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma t)}\,.$
@@
Let $\mathrm{Re}(\gamma)\neq 0$ and also $w\neq 0$. Then (6.24) yields the general solution of the linear system for $\chi_{1}$,
$\chi_{1}=e^{\frac{1}{2}\mathrm{i}\,\varphi}\,\big(c_{1}\,e^{\rho}+c_{2}\,e^{-\rho}\big)\,$,
@@
The above expressions also cover the alternative case $c_{1}=0$ (and $c_{2}\neq 0$) if we allow for $w$ also the negative of the square root in (6.25). If $\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)=0$, then both components of $u^{\prime}$ are functions of $t$ only, multiplied by a phase factor with an exponent linear in $x$. See Appendices Appendix B: Special examples from the class of solutions in Example 6.5 and Appendix C: An almost trivial class of solutions of the two-component vector FL equation for this case. If $\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)\neq 0$, then the wave, described by $u^{\prime}$, travels with constant velocity
@@
As a consequence, $u_{1}^{\prime}$ represents a dark soliton if $\alpha\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\gamma\,(w^{\ast}-1))<0$. If the maximum of the absolute value in the expression for $|u_{1}^{\prime}|^{2}$ is smaller than 1, we have a single dark soliton. An example is shown in Fig. 4.
Figure 4: Plots of the absolute values of the first, and the real part (the imaginary part is similar) of the second component for an $n=1$ solution from the class in Example 6.5. The parameters are $\alpha=\|A\|=2$, $\gamma=1+\mathrm{i}$, $c_{0}=c_{1}=1$, $c_{2}=0$, but we replaced $w$ by w-w. The second plot shows the “beating”. The velocity is $v\approx 0.3$, the beating period $T^{\prime}\approx 1$.
@@
In each case, $u_{2}^{\prime}$ describes a bright soliton (on vanishing background). Solutions of the kind described here already appeared in [7] (see Section 5.2 therein). $\Box$
@@
Let $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$. Then the two roots of the characteristic equation (6.13) are
@@
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,e^{r(x-\frac{k}{\alpha}\,t)}\,c_{1}\,,\qquad\chi_{2}=0\,$,
where $c_{1}\neq 0$ and $r$ now stands for any of the roots $r_{\pm}$. As a consequence of $\chi_{2}=0$, the resulting solution is equivalent to a solution of the scalar FL equation. The Lyapunov equation reduces to the constraint
$k\,\big|r-\frac{\mathrm{i}}{2}(\alpha+\frac{1}{k})\big|^{2}=\alpha^{2}\|A\|^{2}\,$,
which requires $k>0$. This is only compatible with the expression for $r$ if it is real, in which case the constraint is automatically satisfied. In the following, we therefore assume that $r$ is real. The differential equations for $\Omega$ read
$\Omega_{x}=\frac{|c_{1}|^{2}}{2\,k^{2}}\big(1+2\,\mathrm{i}\,k\,(r+\mathrm{i}\,\frac{\alpha}{2})\big)\,e^{2\theta}\,,\qquad\Omega_{t}=-\frac{k}{\alpha}\,\Omega_{x}\,$,
@@
$\theta=r\,\big(x-\frac{k}{\alpha}\,t\big)\,.$
(i) $r\neq 0$. We obtain
$\Omega=\frac{\mathrm{i}\,|c_{1}|^{2}\,(1+\mathrm{i}\,k\,(2r+\mathrm{i}\,\alpha))}{2\,k^{2}\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))}\,e^{2\theta}+c\,$,
@@
$u_{\mu}^{\prime}=A_{\mu}\,e^{\mathrm{i}\,\varphi}\,\Big(1+\frac{k\,|c_{1}|^{2}\big(k\,(2r+\mathrm{i}\,\alpha)-\mathrm{i}\,(1+2\,\alpha\,k)\big)\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))\,e^{2\theta}}{\alpha\|A\|^{2}\big(\mathrm{i}\,|c_{1}|^{2}(1+\mathrm{i}\,k\,(2r+\mathrm{i}\,\alpha))\,e^{2\theta}+2\,k^{2}\,c\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))\big)}\Big)\,,\quad\mu=1,2\,.$
(ii) $r=0$. Then we have
$k=2\,\|A\|^{2}-\alpha^{-1}\pm 2\sqrt{\|A\|^{2}(\|A\|^{2}-\alpha^{-1})}\,$,
which is real if and only if $\|A\|^{2}\geq\alpha^{-1}$. In case of the plus sign, $k$ is then positive. In case of the minus sign, $k$ is only positive if $\alpha<0$. Since now $\theta=0$, we obtain
@@
$u^{\prime}$ $=$ $e^{\mathrm{i}\,\varphi}\,\Big(1-\frac{\mathrm{i}\,|c_{1}|^{2}}{2\alpha\,\|A\|^{2}\,\Omega}\big(\alpha+\frac{1}{k}\big)\Big)\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)$ (6.32)
$=$ $e^{\mathrm{i}\,\varphi}\,\Big(1-\frac{\mathrm{i}\,k\,(1+k\alpha)}{\|A\|^{2}\,\big((1-k\alpha)\,(\alpha\,x-k\,t)+\mathrm{i}\,k\alpha\big)}\Big)\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)\,$, (6.35)
which represents a soliton rationally concentrated along the line $x=(k/\alpha)\,t$ (hence with velocity $k/\alpha$), where $|u^{\prime}_{\mu}|=|A_{\mu}|\,|1-(\alpha^{-1}+k)/\|A\|^{2}|$. This shows that we also obtain dark solitons in this way. An example is displayed in Fig. 6.777This solution is in the “repeated real root” class, obtained via the “limit technique” in [7], page 206.
@@
$\Box$
In case of a double root, i.e., $r_{+}=r_{-}$, (6.24) does not constitute the complete solution of the corresponding part of the linear system. Then we have to find the missing solution, which is done in the next example. This leads, in particular, to quasi-rational solutions.
@@
We have a double root if and only if $r=0$, so that